OpenStudy (anonymous):

f(x)=8+8(6-x)^0.5 / f(x)=x+2 I came up with: 8+8(4-x)^0.5 with a domain of x<=4 Did I do it right?

OpenStudy (anonymous):

Is this the equation that you are working with? \[\frac {8+8*(6-x)^{0.5}}{x+2}\]

OpenStudy (anonymous):

let me know if what i wrote is correct or not?

OpenStudy (anonymous):

I understand why I confused you. Here's what need help with. f(g(x)) where f(x)=8+8sqrt{6-x); g(x)=x+2

OpenStudy (anonymous):

okay, then your answer looks right

OpenStudy (anonymous):

all of it

OpenStudy (anonymous):

thanks